OpenStudy (anonymous):

what is 5.75 in fraction form

OpenStudy (carniel):

5 and 3/4

OpenStudy (anonymous):

yes it is 5 3/4

OpenStudy (anonymous):

575/100

OpenStudy (yanasidlinskiy):

@Carniel was right!!! It is 5&3/4

OpenStudy (anonymous):

Simply, 5 3/4

OpenStudy (anonymous):

@hello1213 Correction: 575/1000

OpenStudy (anonymous):

nope 575/100

OpenStudy (anonymous):

575/100=5.75

OpenStudy (mathmale):

I'd argue that what is 5.75 should be expressed as a single improper fraction. 1. Write 5.75 as 5.75/1. 2. MULTIPLY numerator and denominator by 100. 3. Reduce the resultant improper fraction. Alternatively, Rewrite 5.75 as 5 3/4 and combine the 5 and the 3/4 into a single improper fraction. Note that 5 3/4 is NOT in the form of a fraction.

OpenStudy (anonymous):

@hello1213 oops, yea your right.

OpenStudy (mathmale):

Applause for @hello1213!!!!

OpenStudy (yanasidlinskiy):

my bad:( you are right!!! @mathmale

OpenStudy (carniel):

Its the same thing.....

OpenStudy (anonymous):

You simply simplify 575/100 . And that is you're answer.

OpenStudy (mathmale):

"It" ??? Again, @hello1213 is correct in saying that 575/100 needs to be simplified.

OpenStudy (carniel):

He said make it into a fraction. What I said is correct.

OpenStudy (anonymous):

this is a fraction\[\frac{ 575 }{ 100 }\]

OpenStudy (anonymous):

this is a "mixed fraction"\[5\frac{ 3 }{ 4 }\]

OpenStudy (carniel):

Same thing..... Even so I can get 23/4

OpenStudy (carniel):

He never said what kind of fraction.

OpenStudy (anonymous):

it says" fraction form"

OpenStudy (carniel):

It's still a fraction

OpenStudy (anonymous):

yes but it is unnecessary to put it in "mixed fraction" form.

OpenStudy (yanasidlinskiy):

omg!!! everybody's going crazy!!! lol!!!

OpenStudy (anonymous):

and harder to simplify.

OpenStudy (carniel):

No it isn't..... I don't care what you and Mathmale say. I got it correct.

OpenStudy (anonymous):

okay @Carniel